Fermi National Laboratory
 

Operation Summary for November 23, 2009 - November 25, 2009

 

Monday Nov. 23

The day shift began with Operations monitoring the Tevatron (TeV) store 7383 (119.78E30) and the Recycler stash (104.0E10), with the Antiproton Source stacking (20.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Switchyard (SY) personnel sent beam to the absorber and then on to the Meson MTest area at 10:30 AM. At 11:19 AM, a Linac RF station (LRF5) tripped off and would not reset. Linac Personnel investigated. Operations loaded an energy conservation timeline into the Time Line Generator (TLG). Recycler personnel conducted studies.

During the evening shift, Linac personnel reported at 7:58 PM that they had completed repairs on LRF5. The following devices caused shot downtimes during the shift: a Linac Klystron RF station (KRF5), LRF5, a P1 transfer line area power supply (HT701) went low and caused a beam permit trip, and a NuMI power supply (V100E) went high and caused a beam permit trip.

Tuesday Nov. 24

The midnight shift began with Operations monitoring store 7383 (17.35E30) and the stash (251.4E10), with Pbar stacking (20.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7383 and put the TeV into shot setup at 12:22 AM. Operations established store 7384 at 1:43 AM with an average initial luminosity (AIL) of 192.2E30. Operations began sending beam to the MTest area at 4:17 AM. At 6:41 AM, the MI-52 kicker magnet current dropped; operators investigated, but could find nothing wrong. Operators put Pbar in standby while EE Support personnel investigated.

During the day shift, EE Support personnel completed the MI-52 kicker high voltage power supply repairs by 9 AM. Operations delivered beam to the MTest area until 2:50 PM. MI and Booster personnel checked proton coalescing parameters and controls personnel performed various ACNET work.

During the evening shift, TeV personnel conducted end-of-store studies from 6 PM to 8 PM. Operations terminated store 7384 at 8:15 PM. Operations put the TeV into shot setup and established store 7385 at 9:45 PM with an AIL 279E30. At 11:37 PM, LRF5 tripped off and would not reset; operators called in a Linac technician.

Wednesday Nov. 25

The midnight shift began with Operations monitoring store 7385 (179.13E30), the stash (99.5E10), and the stack (12.5E10), and with no beam to Booster, MI, Pbar, and the NuMI and MiniBooNE experiments due to problems with LRF5.

During the midnight shift, the Linac tech brought LRF5 online at 12:27 AM. MIRF15 tripped off and would not reset; operators put the station into standby.

The Plans for Wednesday and the Holiday Weekend

The plans for today and the long holiday weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/25/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory